OpenStudy (anonymous):

. Solve 4 – 2q = -5(q + 1). Please show all of your work. 4+2q+5q-4=-5q+5q+ 5-4 7q=1 7q/7=1/7 q= 1/7 What did I do wrong?

OpenStudy (anonymous):

you changed the sign of 2q

OpenStudy (anonymous):

q should equal -3

OpenStudy (anonymous):

andd you did that^

OpenStudy (anonymous):

also you distributed your -5 incorrectly

OpenStudy (anonymous):

4 - 2q = -5q-5 9 -2q = -5q 9 = -3q q = 9/-3 q = -3

OpenStudy (anonymous):

can u show me how you would of done the problem 4 – 2q = -5(q + 1)

OpenStudy (anonymous):

4-2q=-5(q+1) distribute the -5 4-2q=-5q-5 move the qs to one side 4=-5q-5+2q move the constants to one side 4+5=-5q+2q simplify: 9=-3q divide by -3 -3=q or q=-3

OpenStudy (anonymous):

ok i see it now. I should of not changed the sign that's how it's wrong. thanks.

OpenStudy (anonymous):

np